OpenStudy (anonymous):

2x sqquare root 2+2x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[2x \sqrt{20+2x}\]

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

\[2x \sqrt{20+2x}\]

OpenStudy (anonymous):

do we solve for x if so the rhs. is missing

OpenStudy (anonymous):

no jusst simplify

OpenStudy (anonymous):

\[2\sqrt{2}x^{3/2} + 2\sqrt{10}x^{1/2}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome (press the good answer button or put an ending post so that other people going through this on the group don't continue to go through it )